### v1

# The Oracle and The Prism: A Decoupled and Efficient Framework for Generative Recommendation Explanation

Abstract.
The integration of Large Language Models (LLMs) into explainable recommendation systems often leads to a performance-efficiency trade-off in end-to-end architectures, where joint optimization of ranking and explanation can result in suboptimal compromises. To resolve this, we propose Prism, a novel decoupled framework that rigorously separates the recommendation process into a dedicated ranking stage and an explanation generation stage.
Inspired by knowledge distillation, Prism leverages a powerful teacher LLM (e.g., FLAN-T5-XXL) as an Oracle to produce high-fidelity explanatory knowledge. A compact, fine-tuned student model (e.g., BART-Base), the Prism, then specializes in synthesizing this knowledge into personalized explanations. This decomposition ensures that each component is optimized for its specific objective, eliminating inherent conflicts in coupled models.
[…]
Recommender Systems, Explainable Recommendation, Large Language Models, Generative Explanation, Knowledge Distillation
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXisbn: 978-1-4503-XXXX-X/18/06conference: Submitted for Review; ; $c$s: Information systems Recommender systemsccs: Computing methodologies Natural language processing
1. Introduction
Recommender systems (lu2012recommender) have become essential in today’s digital landscape (deldjoo2024review), helping users navigate vast information spaces (ricci2011recommender). However, the growing complexity of these systems, particularly with deep learning architectures (he2017neural), creates a “black-box” problem (hassija2024interpreting; csahin2025unlocking) that undermines user trust (zhang2020explainable_survey). Explainable Recommendation, a key area within Explainable AI (XAI(gunning2019xai), addresses this challenge by providing transparent justifications for recommendations. High-quality explanations not only enhance system transparency but also increase persuasiveness, foster user trust, and support better decision-making (lubos2024llm; vultureanu2022survey). Despite these benefits, generating explanations that are both faithful to the model’s reasoning and naturally personalized remains a significant challenge.
Early explainable recommendation methods, such as revealing knowledge graph paths (wang2019knowledge) or influential neighbors in collaborative filtering (sarwar2001item), offered limited transparency and lacked natural language fluency. The rise of Large Language Models (LLMs) has transformed the field, enabling more natural and personalized explanations (fan2023recommender; wu2023survey). Works like XRec (ma2024xrec) propose end-to-end frameworks that jointly optimize recommendation and explanation generation. However, ranking accuracy and explanation quality are not always aligned: coupled models may favor easy-to-explain items at the expense of recommendation performance, or produce hallucinated explanations that misrepresent the true reasoning behind recommendations.
To address these limitations, we propose Prism, a novel decoupled framework for generative explanation in recommender systems. Inspired by augmentation-based paradigms like KAR (xi2024towards), which successfully separate LLM-based reasoning from traditional ranking, we extend this decoupling philosophy to explanation generation. Our framework consists of two independent stages: the Ranking Stage employs any state-of-the-art recommender to determine what to recommend, while the Explanation Stage utilizes our fine-tuned Prism model to generate why it was recommended.
The development of Prism is based on a knowledge distillation pipeline (hinton2015distilling), where we leverage a powerful teacher LLM (FLAN-T5-XXL) to create a large-scale instruction-tuning dataset (wei2021finetuned; ouyang2022training). To meet task-specific interpretability requirements, we adapt the generative paradigm of GenRec (ji2024genrec)—originally designed for recommendation—to fine-tune a compact student model (BART-Base) specifically for explanation generation. By integrating user-aware information through GenRec’s architecture, Prism produces highly personalized explanations.Unlike KAR (xi2024towards), which employs “LLM-assisted ranking,” Prism is the first framework to achieve a complete decoupling between ranking and explanation generation — the output of the ranking stage is used solely as the input condition for the explanation stage, with no joint training or parameter sharing. This design enables Prism to plug into any recommender system (e.g., Collaborative Filtering (sarwar2001), KGCN (wang2019), Deep Interest Network), breaking free from the dependency of coupled frameworks on a single model.
Our main contributions are summarized as follows:
  • We propose Prism, a fully decoupled generative framework that rigorously separates ranking and explanation tasks. This design directly resolves the objective conflict inherent in coupled models, allowing each component to specialize without compromise.
  • We introduce a faithfulness-constrained knowledge distillation approach to ensure semantic alignment between explanations and the user’s interaction history, avoiding the need for complex joint training.
  • We demonstrate that distillation not only transfers knowledge but also serves as a noise filter. The student model (Prism) corrects factual hallucinations present in the teacher’s outputs, leading to more robust and faithful explanations.
2. Related Work
Our work bridges Explainable Recommender Systems and the application of Large Language Models (LLMs) in recommendation (wang2024towards). We review relevant literature to contextualize our contribution.
2.1. Explainable Recommender Systems
Explainable Recommendation has long sought to enhance the transparency of recommender systems (zhang2020explainable_survey). Traditional methods include:
  • Neighborhood-based methods (e.g., Item-based Collaborative Filtering (sarwar2001item)) explain recommendations by showing similar items or users. While intuitive, they rely solely on collaborative signals.
  • Matrix factorization-based methods attempt to interpret latent factors, though these often lack clear semantics.
  • Knowledge Graph-based (KG-based) methods (wang2019knowledge) provide structured explanations via paths in a knowledge graph, offering better interpretability.
Despite their contributions, these approaches typically produce rigid, template-based explanations that lack the fluency and personalization of natural language.
2.2. Large Language Models for Recommendation
The emergence of LLMs has introduced new paradigms for recommendation (fan2023recommender; wu2023survey), which can be categorized by their degree of coupling:
  • Augmentation-based Paradigm: This soft decoupling uses LLMs as external knowledge reasoners. For example, KAR (xi2024towards) employs an LLM to infer textual knowledge for augmenting a traditional ranker’s features. While the LLM assists the ranking process, the final recommendation still depends on the traditional model.
  • Coupled Paradigm: This end-to-end approach uses a single LLM for both understanding and ranking. GenRec reframes recommendation as a sequence generation task, fine-tuning an LLM to directly generate item titles. Although elegant, this requires the LLM to learn complex collaborative patterns from scratch.
Our work, Prism, introduces a third paradigm: a fully decoupled, generative framework. Unlike KAR (where the LLM enhances the ranker) and GenRec (where the LLM acts as the ranker), Prism treats the ranking model as a black-box item selector and employs a specialized LLM solely for explanation generation. This strict separation allows each component to excel independently, avoiding compromises between accuracy and explainability.
2.3. LLM-based Explanation Generation
Using LLMs for natural language explanations represents a major advance in explainable AI. Current state-of-the-art approaches primarily use coupled, multi-task frameworks. For instance, XRec employs a unified model that jointly learns recommendation and explanation generation. While aiming for consistency, this coupling often forces a trade-off between ranking accuracy and explanation quality.
In contrast, Prism explores a decoupled framework. To our knowledge, it is the first to adapt a generative recommendation architecture (GenRec) specifically for explanation generation within a fully decoupled system. Instead of joint training, we ensure alignment through knowledge distillation, where a teacher model generates faithful explanations to train a smaller student model, enabling specialized optimization for each task.
3. Preliminaries
In this section, we formally define the explanation generation task and outline the sequence-to-sequence (Seq2Seq) architecture (karatzoglou2018seq2seq) that underpins our framework.
3.1. Problem Formulation
Let $\mathcal{U}$ denote the set of users and $\mathcal{I}$ the set of items. Each user $u\in\mathcal{U}$ is associated with a chronological interaction history $H_{u}=(i_{1},i_{2},\dots,i_{t})$, where $i_{k}\in\mathcal{I}$. Given a recommendation pair $(u,i_{\text{rec}})$, where $i_{\text{rec}}\in\mathcal{I}$ is recommended to $u$, our goal is to generate a personalized, faithful, natural language explanation $E=(y_{1},y_{2},\dots,y_{n})$, with tokens $y_{k}$ drawn from a vocabulary $\mathcal{V}$.
We learn a parameterized function $f_{\theta}$ that models the conditional probability:
(1) $P(E\mid H_{u},i_{\text{rec}})=\prod_{k=1}^{n}P(y_{k}\mid y_{<k},H_{u},i_{\text{rec}};\theta)$
Our framework, Prism, optimizes $\theta$ to maximize faithfulness and personalization in generated explanations.
3.2. Sequence-to-Sequence Models for Generation
We build on a pretrained Seq2Seq model—specifically the BART architecture (lewis2020bart)—comprising an Encoder and a Decoder.
The encoder maps an input sequence $X=(x_{1},\dots,x_{m})$ to contextualized hidden states $\mathbf{h}=(\mathbf{h}_{1},\dots,\mathbf{h}_{m})$, capturing the full input context.
The decoder generates $E=(y_{1},\dots,y_{n})$ in an autoregressive manner, predicting each token:
(2) $y_{k}\sim P(y\mid y_{<k},\mathbf{h};\theta)$
Conditioned on $\mathbf{h}$ and past outputs, the model captures user–item context for explanation generation. We train the entire model end-to-end via cross-entropy loss between predicted and reference tokens.
4. Methodology
In this section, we present our proposed Prism, a decoupled framework for generative explanation in recommender systems. Our approach is designed to synergize the ranking strengths of specialized recommendation models with the nuanced reasoning and generation capabilities of Large Language Models (LLMs). As illustrated in Figure 1, the framework is comprised of two distinct stages: an offline stage for creating a high-quality dataset and fine-tuning our explanation model, and an online stage where the model serves as a plug-in explanation module.
Figure 1. The overall framework of Prism. The offline stage consists of a teacher phase for data creation via knowledge distillation and a student phase for model fine-tuning. The online stage demonstrates how Prism functions as a decoupled module alongside any SOTA recommender.
4.1. Overall Framework
The primary focus of this research is not to propose a novel decoupling mechanism or user embedding algorithm. Instead, it aims to, for the first time, systematically investigate the feasibility and effectiveness of successfully and creatively adapting an existing, complex generative framework designed for recommendation (GenRec) to a fully decoupled, downstream explanation generation task via knowledge distillation.
The Offline Stage is where our explanation model is developed:
  • Teacher Phase (Knowledge Distillation): We employ a powerful, large-scale teacher LLM to generate high-quality, ”golden” explanations for given user-item interactions. This process is detailed in Section 4.2.
  • Student Phase (Model Fine-tuning): We then use this distilled dataset to fine-tune a much smaller, more efficient student LLM, adapting it to become a specialist in generating personalized recommendation explanations. This is detailed in Section 4.3.
The Online Stage represents the deployment scenario. Our trained Prism model operates as an independent module, receiving the output from any primary SOTA ranking model and generating a natural language explanation in real-time.
4.2. Knowledge Distillation for Data Creation
A major bottleneck for training high-quality explanation models is the limited availability of large-scale, human-annotated datasets. To address this, we adopt knowledge distillation (hinton2015distilling), using a powerful teacher LLM $\mathcal{M}_{teacher}$ to automatically construct our training corpus.
Teacher Model. We select FLAN-T5-XXL (1111B parameters) for its strong instruction-following and reasoning ability.
Faithfulness-Constrained Prompting. The quality of the distilled dataset depends critically on the prompt guiding the teacher. To reduce factual hallucinations, we design a constraint-driven template explicitly instructing the model to base explanations solely on the user’s interaction history (tonmoy2024comprehensive; lyu2023faithful):
Generate a short, personalized, and persuasive explanation for the following recommendation.
Context:
- User’s movie viewing history: {history}
- Recommended movie: {item_to_explain}
Task: Explain WHY this is a good recommendation based on the user’s history.
- Be specific: Link features of the recommended movie (e.g., genre, director, actors, theme) to patterns in the history.
- Be natural: Sound like a genuine recommendation from a friend.
- Be concise: Ideally one or two sentences.
- Start the explanation directly.
Explanation:
For each raw sample $(H_{u},i_{rec})$ we format the prompt $X_{prompt}$ and obtain the golden explanation:
(3) $E_{golden}=\mathcal{M}_{teacher}(X_{prompt})$
Repeating this over the entire dataset yields the instruction-tuning set:
(4) $\mathcal{D}=\{(X_{j},u_{j},E_{j})\}_{j=1}^{N}$
where $X_{j}$ is the prompt text, $u_{j}$ the user ID, and $E_{j}$ the golden explanation.
Algorithm 1 Knowledge Distillation Pipeline for Explanation Dataset Creation
1:Input: Raw logs $\mathcal{D}_{raw}=\{(u,H_{u},i_{rec})\}$, Teacher LLM $\mathcal{M}_{teacher}$, Prompt Template $T_{prompt}$
2:Output: Explanation dataset $\mathcal{D}_{exp}$
[…]
4:for each $(u,H_{u},i_{rec})$ in $\mathcal{D}_{raw}$ do
5:  $X_{prompt}\leftarrow\mathrm{format}(T_{prompt},H_{u},i_{rec})$
6:  $E_{golden}\leftarrow\mathcal{M}_{teacher}(X_{prompt})$
7:  if $E_{golden}$ is not an error then
8:   Append $(u,H_{u},i_{rec},E_{golden})$ to $\mathcal{D}_{exp}$
9:  end if
10:end for
11:return $\mathcal{D}_{exp}$
4.3. Explanation Model Fine-tuning
Student Model. We choose BART-Base (140M parameters) as our student model. This choice is motivated by its strong performance as an Encoder-Decoder model and its native compatibility with the underlying architecture of the GenRec framework, which we adapt for our task.
User-Aware Input Representation. A key aspect of our approach is to make the explanation model user-aware. We achieve this by adapting to the modified BART architecture within the GenRec framework, which includes a dedicated user embedding layer. Let $W_{u}\in\mathbb{R}^{|\mathcal{U}|\times D}$ be the user embedding matrix, where $|\mathcal{U}|$ is the total number of users and $D$ is the hidden dimension size of the model. This embedding matrix, $W_{u}$, is **randomly initialized** at the beginning of the training process.
Given an input token sequence $X=(x_{1},\dots,x_{m})$, the model first projects each token into a vector space using the standard word embedding matrix $W_{e}$, resulting in $H^{(0)}=(e_{1},\dots,e_{m})$, where $e_{j}=W_{e}(x_{j})$. For the corresponding user $u$, we retrieve their unique user vector $v_{u}=W_{u}(u_{id})$. This user vector is then added to each word embedding in the sequence:
(5) $\hat{e}_{j}=e_{j}+v_{u}$
The final, user-aware input representation for the encoder is thus $\hat{H}^{(0)}=(\hat{e}_{1},\dots,\hat{e}_{m})$. Crucially, during the fine-tuning process, the user embedding matrix $W_{u}$ is trained jointly with all other parameters of the BART model (including $W_{e}$ and the Transformer layers). The gradients from the cross-entropy loss (Equation 4) are backpropagated through the entire model, allowing $W_{u}$ to learn meaningful, user-specific representations that are beneficial for the explanation generation task.
Objective Function. The fine-tuning process aims to minimize the standard cross-entropy loss over the distilled dataset $\mathcal{D}$. Let $E=(y_{1},\dots,y_{n})$ be the sequence of tokens in a golden explanation. The loss for a single sample $(X,u,E)$ is the negative log-likelihood:
(6) $\mathcal{L}(\theta)=-\sum_{t=1}^{n}\log P(y_{t}|y_{<t},X,u;\theta)$
where the probability is now also conditioned on the user $u$. This loss is optimized over the entire training dataset using the AdamW optimizer.
4.4. Scalability and Cold-Start Handling
A potential concern with any user-embedding-based approach is its scalability to millions of users and its performance in cold-start scenarios where new users have no historical data to train their embeddings. Our proposed decoupled framework, however, is inherently robust to these challenges.
The primary responsibility of handling cold-start recommendation lies with the upstream Ranking Module. This module is treated as a black box in our framework and can employ its own specialized strategies (e.g., content-based filtering, contextual bandits) to generate a relevant recommendation for new users. Our Explanation Module (Prism) only activates after a recommendation has already been successfully made.
In a cold-start scenario where a user_id is new and has no trained embedding in our $W_{u}$ matrix, our framework can gracefully handle the situation by assigning a default “unknown user” embedding (e.g., a zero vector). In this case, the user-aware component is effectively disabled, causing Equation(̃5) to simplify to $\hat{e}_{j}=e_{j}$. The model then defaults to generating a high-quality, but non-personalized, content-based explanation. Crucially, it still produces a relevant explanation because its primary conditioning signal is the rich textual information from the user’s (potentially short) history $H_{u}$ and the recommended item $i_{rec}$, not the user ID itself.
Therefore, unlike monolithic models where an unknown user embedding might cripple the entire recommendation process, our decoupled design ensures that the system never fails. It simply degrades gracefully from a ”personalized explainer” to a still highly effective ”content-based explainer” in the face of unknown users. This robustness is a key architectural advantage of our approach.
5. Experiments
In this section, we detail the experimental setup designed to rigorously evaluate our proposed Prism framework.
5.1. Research Questions (RQs)
Our experiments are designed to answer the following key research questions:
  • RQ1 (Overall Performance): Can our fine-tuned Prism model generate higher quality explanations than a strong, zero-shot large language model baseline in terms of both automatic and human-evaluated metrics?
  • RQ2 (Ablation Study): Does the user-aware mechanism, adapted from the GenRec framework, demonstrably contribute to the personalization of the generated explanations?
  • RQ3 (Qualitative Analysis & Robustness): What are the qualitative characteristics of the explanations generated by Prism? Specifically, does our framework exhibit robustness against the factual hallucinations present in the teacher model’s distilled knowledge?
5.2. Dataset
To assess the performance and generalization of our framework, we experiment on two widely used public benchmarks: MovieLens-1M (tousch2019robust; harper2015movielens) and Yelp (luca2016fake).
MovieLens-1M contains $\sim$1 million explicit ratings from 6,040 users on 3,883 movies and is a standard benchmark in recommender system research. Yelp presents a more diverse and realistic scenario, comprising user reviews of local businesses across multiple categories, thereby capturing a broad range of real-world preferences.
We preprocess both datasets by converting raw user interactions into chronological sequences and truncating each user’s history to their most recent 50 interactions. This choice balances computational efficiency with sufficient behavioral context and aligns with the empirical distribution—over $95\%$ of MovieLens-1M users have sequences of length $\leq 50$. This design ensures transformer models receive representative input lengths without excessive overhead. The effect of history length on performance remains an interesting topic for future study.
Final preprocessed dataset statistics are reported in Table 1, and all evaluations are conducted on the full test sets of both domains.
Table 1. Statistics of the processed datasets.
Statistic MovieLens-1M Yelp
# Users 6,040 1,987,929
# Items 3,883 150,346
# Train Sequences 894,752 1,418,452
# Test Sequences 99,417 157,606
5.3. Baselines
We evaluate our proposed Prism framework against a comprehensive suite of baselines, covering classic, recent state-of-the-art, and large-scale zero-shot models.
  • Att2Seq (dong2017att2seq): A classic and strong baseline from the pre-LLM era. It utilizes an attention-based sequence-to-sequence GRU model to generate textual outputs, allowing us to measure the performance leap brought by modern pre-trained transformer architectures.
  • PEPLER (li2023personalized_pepler): An advanced framework that leverages a PE-enhanced PLM for explanation generation, representing a strong recent baseline.
  • FLAN-T5-XXL (Zero-Shot): This 11B parameter teacher model represents the upper-bound performance of a massive, general-purpose LLM on our task without any domain-specific fine-tuning.
  • BART-Base (Zero-Shot): This is the same 140M parameter base architecture as our Prism model. This baseline is crucial for isolating the performance gains attributable solely to our knowledge distillation and fine-tuning pipeline.
5.4. Comparison with Coupled Frameworks
A note on comparison with coupled frameworks (e.g., XRec): Our empirical focus is on validating the decoupled paradigm itself. A direct, apples-to-apples comparison with a coupled model is challenging due to fundamental differences in architecture and objective functions. Coupled models are trained to optimize a joint loss, and their explanation quality is tied to their internal ranking mechanism. In contrast, Prism is designed to explain the output of any black-box ranker. Therefore, we position Prism as a complementary, alternative paradigm suited for scenarios prioritizing modularity and efficiency, and we compare it against strong generative and zero-shot baselines to isolate the efficacy of our distillation approach for explanation generation.
5.5. Evaluation Metrics
To comprehensively evaluate our approach, we employ both automatic and human evaluation protocols.
Automatic Evaluation. We adopt the following established metrics:
  • ROUGE (lin2004rouge): Measures n-gram overlap between generated and reference texts. We report F1 scores for ROUGE-1 (unigram), ROUGE-2 (bigram), and ROUGE-L (longest common subsequence). ROUGE-N is computed as:
    (7) $\text{ROUGE-N}=\frac{2\cdot P_{n}\cdot R_{n}}{P_{n}+R_{n}}$
    where $P_{n}$ and $R_{n}$ denote n-gram precision and recall.
  • BERTScore (zhang2019bertscore): Computes semantic similarity using contextual embeddings from RoBERTa-Large. For prediction $p_{i}$ and reference $r_{j}$, similarity is measured via cosine similarity $x_{i}^{T}x_{j}$. We report the F1 variant:
    (8) $F1_{\text{BERT}}=\frac{2\cdot P_{\text{BERT}}\cdot R_{\text{BERT}}}{P_{\text{BERT}}+R_{\text{BERT}}}$
  • GPTScore (fu2023gptscore): Assesses fluency and coherence using a generative LLM as evaluator. The score is the length-normalized log-likelihood of explanation $E=(y_{1},\dots,y_{n})$ given context $C$:
    (9) $\text{GPTScore}(E,C)=\frac{1}{n}\sum_{k=1}^{n}\log P_{\mathcal{M}_{\text{eval}}}(y_{k}\mid y_{<k},C)$
    where $\mathcal{M}_{\text{eval}}$ is GPT-3.5-Turbo. Higher scores indicate more natural explanations.
Human Evaluation. We conduct a user study with 30 graduate students, who rate explanations on a 5-point Likert scale for:
  • Persuasiveness: Likelihood of convincing the user to watch the movie
  • Personalization: Degree of tailoring to the user’s history
  • Faithfulness: Factual grounding in the user’s history
Inter-annotator agreement (Fleiss’ Kappa) is reported to ensure reliability.
5.6. Efficiency Analysis
Beyond explanation quality, computational efficiency is crucial for real-world deployment (roy2022systematic). We evaluated the practical viability of our framework by measuring inference latency and peak GPU memory usage for our fine-tuned Prism model against the massive FLAN-T5-XXL baseline.
[…]
These gains validate our knowledge distillation approach: we successfully compress and transfer explanatory knowledge from a large, expensive, hard-to-deploy teacher model into a compact, fast, and deployable student, without notable loss in human-perceived quality (cf. human evaluation). This demonstrates that our decoupled framework offers a practical and cost-efficient solution for high-quality explainable recommendation in production environments.
Table 2. Efficiency comparison. Latency is the average time to generate one explanation over 100 runs.
Model Params Latency (ms) Peak GPU (GB)
FLAN-T5-XXL 11B 4612.92 20.60
Prism 140M 190.30 1.91
[…]
6. Results and Analysis
In this section, we present and analyze the empirical results of our experiments. We aim to answer our research questions by quantitatively comparing our model against the baseline, conducting a targeted ablation study, and performing an in-depth qualitative analysis.
6.1. Overall Performance (RQ1)
To answer our first research question, we conducted a comprehensive evaluation on two distinct datasets. The main experimental results, encompassing both automatic and human evaluations, are presented in Table 3.
Table 3. Main experimental results on both Yelp and MovieLens-1M datasets. For human evaluation, we report mean score $\pm$ standard deviation. IAA using Fleiss’ Kappa was 0.75.
Dataset Model Automatic Metrics Human Evaluation
ROUGE-L GPTScore BS-F1 Persuasive. Personal. Faithful.
Yelp FLAN-T5-XXL (11B) 0.2761 68.12 0.2679 3.01 $\pm$ 0.88 3.07 $\pm$ 0.92 2.87 $\pm$ 1.01
Att2Seq 0.1653 63.91 0.2377 2.98 $\pm$ 0.85 2.76 $\pm$ 0.90 2.95 $\pm$ 0.98
BART-Base (140M) 0.1607 62.59 0.2599 3.05 $\pm$ 0.96 3.13 $\pm$ 0.81 3.05 $\pm$ 0.95
PEPLER 0.2002 67.24 0.3032 3.42 $\pm$ 0.85 3.35 $\pm$ 0.77 3.11 $\pm$ 0.91
Prism (w/o User) 0.2183 69.21 0.3273 3.85 $\pm$ 0.65 3.62 $\pm$ 0.80 3.87 $\pm$ 0.69
Prism (Full) 0.2259 71.56 0.3334 3.99 $\pm$ 0.63 4.02 $\pm$ 0.67 4.06 $\pm$ 0.65
MovieLens-1M FLAN-T5-XXL (11B) 0.2953 74.09 0.2886 3.31 $\pm$ 0.82 3.29 $\pm$ 0.96 2.92 $\pm$ 0.89
Att2Seq 0.1712 70.97 0.2591 3.14 $\pm$ 0.84 3.02 $\pm$ 0.80 2.91 $\pm$ 0.95
BART-Base (140M) 0.1764 71.33 0.2690 3.03 $\pm$ 0.75 3.11 $\pm$ 0.77 3.20 $\pm$ 0.88
PEPLER 0.2130 73.01 0.3246 3.35 $\pm$ 0.72 3.40 $\pm$ 0.74 3.36 $\pm$ 0.88
Prism (w/o User) 0.2478 79.52 0.3545 3.90 $\pm$ 0.68 3.63 $\pm$ 0.85 4.05 $\pm$ 0.62
Prism (Full) 0.2574 80.74 0.3589 4.03 $\pm$ 0.61 4.07 $\pm$ 0.59 4.12 $\pm$ 0.57
Analysis of Results. The comprehensive results in Table 3 lead to several key conclusions.
First, our proposed Prism (Full) model consistently and significantly outperforms all baselines in the crucial human evaluation metrics across both datasets. For instance, on the MovieLens-1M dataset, its Faithfulness score of 4.12 is substantially higher than the strongest baseline, PETER (3.36), and the massive FLAN-T5-XXL (2.92). This trend holds on the more challenging Yelp dataset, validating that our decoupled knowledge distillation framework successfully trains a student model that generates explanations perceived by humans as more persuasive, personalized, and trustworthy.
Second, the automatic metrics reveal a more nuanced story. On metrics that measure semantic similarity like GPTScore and BERTScore-F1, our Prism models also achieve state-of-the-art performance, surpassing the 11B parameter FLAN-T5-XXL. This suggests our fine-tuned model better captures the semantic essence of a good explanation. However, on the lexical overlap metric ROUGE-L, the zero-shot FLAN-T5-XXL baseline achieves the highest score. This finding supports our hypothesis that a high ROUGE score can be misleading, as it rewards the stylistic self-consistency of the teacher model’s outputs—which, as our qualitative analysis shows, often contain factual hallucinations.
Comparison with Coupled Frameworks. In this study, we focus our empirical comparison on generative and zero-shot baselines, rather than directly including coupled frameworks.
Figure 2. Automatic evaluation results on ROUGE-L, GPTScore, and BS-F1 metrics across Yelp and MovieLens-1M datasets.
Figure 3. Human evaluation results on persuasiveness, personalization, and faithfulness dimensions.
This is a deliberate methodological choice aligned with our primary research objective: to validate the viability of a fully decoupled paradigm. A direct comparison between our specialized explanation generator and a monolithic model trained for both ranking and explanation would be challenging to interpret, as their fundamental objectives and architectural assumptions differ significantly. Our selected baselines enable a controlled evaluation of our core contribution—the efficacy of knowledge distillation for the explanation generation task. We acknowledge that a comparative study between optimized implementations of both paradigms represents an important direction for future work.
6.2. The Pitfalls of Automatic Metrics: A Deeper Look
While Table 3 provides a preliminary performance overview, it also reveals a counter-intuitive phenomenon: the large FLAN-T5-XXL baseline attains the highest ROUGE scores. We argue this is misleading and highlights a major pitfall of relying solely on lexical-overlap metrics for this nuanced task (schluter2017limits).
This inflated ROUGE largely stems from comparing the baseline against its own generated “golden” explanations, rewarding stylistic similarity over factual accuracy. Table 4 illustrates the issue: both the golden explanation (A) and baseline output (B) share entities and phrasing yet contain severe hallucinations—yielding a high ROUGE-L. In contrast, a factually correct but lexically different explanation (C) receives an unfairly low score.Although ROUGE-L has known limitations for this task, it remains a common benchmark in text generation.
Table 4. High lexical overlap between two incorrect statements (A vs. B) results in a higher ROUGE-L than between a correct and incorrect statement (A vs. C).
Context User History: … E.T., Star Wars …, Recommended Item: Back to the Future
A: Golden Explanation Back to the Future is based on ”The Wizard of Oz” and influenced by ”The Phantom Menace”.
B: FLAN-T5 Prediction Back to the Future is a sci-fi movie influenced by ”The Phantom Menace” and ”The Wizard of Oz”.
C: A Faithful Explanation It’s a classic 80s sci-fi adventure, similar to other films in your history.
ROUGE-L (A vs. B) 0.75 (Deceptively High)
ROUGE-L (A vs. C) 0.15 (Unfairly Low)
This case shows that high ROUGE can mask unfaithful explanations, reinforcing the necessity of human evaluation for measuring Faithfulness and Personalization—critical qualities for explainable recommender systems.
Human Evaluation. Automatic metrics like ROUGE, based on lexical overlap, cannot distinguish a factually correct explanation from a fluent hallucination. They reward stylistic similarity even when semantic fidelity is flawed. We therefore treat human judgment as the ultimate ground truth.
We conducted a human study in which annotators scored outputs from all models on Persuasiveness, Personalization, and Faithfulness. Results (Table 3) reveal a fundamentally different picture: Prism is overwhelmingly preferred, significantly outperforming FLAN-T5-XXL and zero-shot BART-Base in all dimensions ($p<0.01$, paired t-test). The largest gains appear in Personalization and Faithfulness, indicating that despite lower ROUGE due to vocabulary differences, Prism effectively learns to produce trustworthy, genuinely helpful explanations—filtering noise from its imperfect teacher.
6.3. Analysis of Teacher Hallucinations and Student Robustness (Addressing RQ3)
To answer RQ3, we conducted a qualitative analysis of the generated explanations. This analysis revealed a remarkable and unexpected phenomenon: our fine-tuned student model, Prism, not only learned to generate fluent explanations but also demonstrated an ability to correct or ignore the factual hallucinations produced by its powerful teacher model.
As shown in the case studies in Appendix B,Table 5, the teacher model (FLAN-T5-XXL) frequently produces non-factual or logically flawed ”hallucinated explanations” (this is inevitable(banerjee2025llms))(marked in red). For instance, it incorrectly associates ”Back to the Future” with ”The Wizard of Oz.” In contrast, our Prism often filters this noise and provides a more conservative but factually correct explanation.
Beyond knowledge transfer, we observed that the fine-tuning process imbues the student model with a degree of robustness against the teacher’s hallucinations. We hypothesize that this stems from a regularization effect inherent in model compression.
6.4. Ablation Studies
The smaller capacity of the 140140M-parameter student model (BART-Base) constrains its ability to fully reproduce the teacher’s output distribution, which contains both valid patterns and occasional errors. Consequently, the student prioritizes salient and coherent patterns from the distilled dataset, implicitly treating extreme hallucinations as outliers. This property suggests our framework serves not only as a distillation method but also as a potential knowledge refinement technique. A deeper investigation (e.g., varying student capacities or architectures) lies beyond this paper’s scope but represents a promising direction for developing more reliable and truthful generative models.
To validate our design choices and understand performance sources, we conduct two ablation studies:
6.4.1. Effectiveness of Knowledge Distillation and Fine-Tuning
We first ask: Is the full knowledge distillation + fine-tuning pipeline necessary? We compare our fully trained Prism with its zero-shot foundation model (BART-Base), which shares the same architecture but has not been fine-tuned on our distilled explanations.
Results in Table 3 reveal a large gap across all metrics. On Yelp, zero-shot BART-Base often produces repetitive or irrelevant content, with BERTScore-F1 of $0.2599$, whereas Prism reaches $0.3334$. This confirms that domain-specific fine-tuning on a high-quality distilled dataset is indispensable for enabling a compact model to handle complex explanation generation.
6.4.2. Impact of the User-Aware Mechanism
We next examine our user-aware input representation (Section 4.3). We train an ablated variant, Prism w/o User, by removing the user-specific embedding, and compare it with Prism (Full Model).
As shown in Table 3, removing the user-aware component causes notable drops, especially in human-evaluated Personalization scores. This empirically confirms that adapting GenRec’s user-aware architecture is a critical factor in generating explanations that feel tailored to individual users.
7. Conclusion and Future Work
This paper addressed the critical challenge of generating high-quality, personalized, and faithful explanations for recommender systems. We identified a fundamental limitation in existing coupled, multi-task frameworks: the inherent trade-off between recommendation accuracy and explanation quality. To overcome this, we introduced Prism, a novel decoupled framework that cleanly separates the ranking and explanation generation tasks. By leveraging knowledge distillation and a user-aware adaptation of the GenRec architecture, Prism demonstrates that a compact, fine-tuned student model can not only compete with but also surpass strong zero-shot baselines and classic attention-based sequence-to-sequence models. Human evaluations particularly highlighted its superiority in terms of persuasiveness, personalization, and faithfulness, with the model even exhibiting a degree of robustness against potential noise from the teacher model. Prism’s lightweight design (140M parameters, $1.91$ GB peak memory) enables edge deployment. In practical e-commerce testing, explanation latency dropped to $190$ ms, meeting real-time Web application requirements. In conclusion, our work provides strong empirical evidence that a decoupled, distillation-based approach is a viable and effective pathway toward building more trustworthy and user-centric recommender systems.
While this study establishes a robust Proof-of-Concept for the decoupling principle, several limitations naturally point to promising avenues for future work:
  • Architectural Exploration: The user-aware mechanism, while effective, is adopted directly from GenRec. Future research could explore more advanced personalization techniques, such as dynamic user embeddings or meta learning strategies for cold-start users(yuan2023user), to further enhance explanation quality.
  • Generalizability and Benchmarking: Our empirical assessment was conducted on two classic datasets, and the scope precluded inclusion of very recent models (e.g., RAG-based) or explicit cross-domain tests. A critical next step is to evaluate Prism against a wider array of state-of-the-art models and across more diverse domains (e.g., e-commerce, news) to thoroughly verify its generalizability.
  • Mechanism Analysis: A fascinating finding is the student model’s emergent capability to filter teacher-generated hallucinations. A deeper dissection of this mechanism through detailed ablation studies (e.g., on model capacity, prompt constraints) and specialized factuality metrics (e.g., FactScore (min2023factscore)) is a pivotal objective.
  • Philosophical Synergy with RAG(arslan2024survey): Our framework shares a philosophical foundation with Retrieval Augmented Generation (RAG). We conceptualize Prism as a form of “Recommendation-Augmented Generation” where the generative model explains a pre-retrieved item. Future work could deepen this connection by having the ranker provide explicit evidence (e.g., key user behaviors) for the recommendation, enabling the generator to produce explanations that are not only faithful but also fully auditable, paving the way for a new generation of transparent recommender systems(fan2024survey).
Acknowledgements.
I am deeply grateful to my supervisor, Prof. Daqiang Zhang. Thank you for trusting a mathematics undergraduate to explore this vast world, and for planting the first seed of academic curiosity in my heart. I also thank the 30 students who helped with the evaluation—your patience was a gift to this project. To my dear friend Fu: though thousands of miles separate Zhuhai from the Northeast, you felt closer than anyone. Thank you for being my emotional safety net. When I was drowning in self-doubt, you didn’t just tell me to swim—you listened to my pain and gently helped me find my way back. To my roommates, thank you for the warmth of your company. Your laughter was the best cure for my anxious days. To my parents, thank you for being my eternal sanctuary. I am especially grateful to my father for your infinite tolerance, which taught me that ”effort is enough” and allowed me to breathe. And to those who know all my imperfections yet still choose to love me—your acceptance gave me the courage to face the world. To the ”me” in those solitary midnights: I know how difficult it was to bridge the gap between mathematics and code. Thank you for that quiet resilience—for breaking down, drying your tears, and always choosing to start again. The world is splendid and vast, and I am but a traveler standing at its gate. May I always remember: never let your soul harden into a mere nail in the machinery, but to remain forever a student of this world—stay hungry, stay foolish.
Appendix A Appendix A Human Evaluation Details
To assess explanation quality beyond textual similarity, we conducted a rigorous and systematic human evaluation study. We recruited 30 graduate students with foundational and advanced knowledge in recommender systems to participate in the evaluation, ensuring domain expertise for accurate and informed judgments.
A.1. Evaluation Procedure
Each annotator was presented with a series of anonymized user historical interactions alongside corresponding generated explanations. For each sample, the annotator rated on a 5-point Likert scale according to:
  • Persuasiveness: Likelihood the explanation convinces the user to watch the movie (1 = Not at all, 5 = Very likely).
  • Personalization: Degree of tailoring to the specific user history (1 = Generic, 5 = Highly personalized).
  • Faithfulness: Factual and logical grounding in user history (1 = Not faithful / Hallucinated, 5 = Very faithful).
A.2. Annotation Guidelines and Training
To ensure consistency and objectivity, detailed guidelines were provided, including:
  1. (1)
    Clear definitions for each dimension.
  2. (2)
    Examples for all score levels (1–5).
  3. (3)
    Instructions to avoid bias by using only the provided history and explanation.
Before the main evaluation, annotators trained on a calibration set of ten samples. Feedback was given, and disagreements resolved to unify scoring standards.
A.3. Independent and Blind Annotation
Annotations were performed independently to avoid influence from other annotators. The annotation interface:
  • Presented user history and explanations clearly.
  • Randomized sample ordering (to avoid position bias).
  • Hid model identity (to avoid source bias).
A.4. Reliability: Fleiss’ Kappa
We calculated Fleiss’ Kappa to measure inter-annotator agreement (IAA) using:
$\kappa=\frac{\bar{P}-\bar{P_{e}}}{1-\bar{P_{e}}}$
where:
$\bar{P}=\frac{1}{N}\sum_{i=1}^{N}P_{i},\quad P_{i}=\frac{1}{n(n-1)}\left[\sum_{j=1}^{k}n_{ij}^{2}-n\right]$
$\bar{P_{e}}=\sum_{j=1}^{k}p_{j}^{2},\quad p_{j}=\frac{1}{Nn}\sum_{i=1}^{N}n_{ij}$
Here, $N$ is the number of items, $n$ the number of annotators, $k$ the number of rating categories, and $n_{ij}$ the number of annotators assigning category $j$ to item $i$. A $\kappa$ above 0.6 indicates substantial agreement (landis1977measurement).
A.5. Statistical Analysis
We computed the mean, median, and standard deviation of ratings for each dimension, and conducted paired t-tests to assess the statistical significance of differences between models.
Appendix B Appendix B Case Study and Analytical Discussion
B.1. Table 5: Case Study of Generated Explanations
This appendix presents a detailed case study (Table 5) derived from the main experiment, illustrating the advantages of the proposed Prism framework in generating personalized, faithful, and persuasive recommendation explanations.
The case compares three models:
  • FLAN-T5-XXL (Teacher) — A powerful large-scale model that, while fluent, tends to produce factual hallucinations.
  • BART-Base (Zero-Shot) — A student model without task-specific fine-tuning, representing a naive baseline.
  • Prism (Ours) — A compact, fine-tuned student model trained via a faithfulness-constrained knowledge distillation pipeline.
As shown in Table 5, Prism consistently avoids hallucinations present in the teacher model, while offering richer personalization than the zero-shot student baseline. This highlights its dual strengths in factual faithfulness and user-tailored content generation.
Table 5. Case study of generated explanations. Our Prism demonstrates higher faithfulness and relevance compared to both the powerful-but-hallucinating teacher model (errors marked in red) and the naive zero-shot student baseline.
User History Recommended Item FLAN-T5-XXL (Teacher) BART-Base (Zero-Shot) Prism (Ours)
… E.T., Star Wars, The Thing … Back to the Future …based on ”The Wizard of Oz’ and influenced by ”The Phantom Menace’. Back to the Future is a great movie. Back to the Future is a science fiction film from 1985, fitting your interest in 80s sci-fi.
… Animal House, Caddyshack … American Pie …similar to The Shining and The Adventures of Pinocchio. This is a comedy that suits you American Pie is a comedy that is similar to the movies you have already seen.
… Ferris Bueller’s Day Off … Sixteen Candles The user has a history of watching comedies, and Sixteen Candles is based on teen angst. Sixteen Candles is a movie. The user has a history of watching movies about high school. Sixteen Candles is based on teen angst.
B.2. Analytical Discussion
As demonstrated in Table 5, the teacher model (FLAN-T5-XXL) often produces hallucinated connections that have no grounding in the provided user history. The zero-shot BART-Base baseline, while free from such hallucinations, generally outputs generic and non-personalized statements.
In contrast, our proposed Prism model generates explanations that are both factually verifiable and deeply personalized, aligning with empirical user preferences.
These qualitative observations reinforce the quantitative results reported in the main paper: Prism outperforms all baselines in Persuasiveness, Personalization, and Faithfulness according to human evaluation. The ability to filter out factual noise from the teacher’s outputs, while enriching personalization, underscores the effectiveness of our faithfulness-constrained distillation approach.